# UVGS: Reimagining Unstructured 3D Gaussian Splatting using UV Mapping

### v1

@@ 1 Spherical Mapping
Spherical Mapping: Spherical mapping [42] is a fundamental technique in computer graphics that is used to project 3D meshes onto a 2D map generally for texture mapping, where a 2D image is wrapped around a 3D object, such as a cylinder or a sphere. However, cylindrical mapping fails to capture the top and bottom parts of the object in the same UV map, and can introduce distortions for objects that extend far in the Z-direction. Hence we opted for spherical mapping the process of which involves converting 3D Cartesian coordinates ($x,y,z$) into spherical coordinates ($\rho,\theta,\phi$) and then mapping these onto a 2D plane. Algorithm [1] explains spherical unwrapping in detail for a single layer(K=1). The same process can be repeated for multiple layers, by keeping a track of opacity values.
@@
1:$3DGS\in\mathbb{R}^{n\times 14}$$(M,N)\in\mathbb{Z},~{}K=1$
2:$position(\sigma),color(c),scale(s)\in\mathbb{R}^{n\times 3}$
3:$rotation(r)\in\mathbb{R}^{n\times 4},~{}opacity(o)\in\mathbb{R}^{n\times 1}$
@@
8:$(\theta,~{}\phi)\leftarrow(\text{deg}(\theta)+180,~{}\text{deg}(\phi))$
9:$\theta_{UV}\leftarrow\text{round}((\theta/360)\times M)$
10:$\phi_{UV}\leftarrow\text{round}((\phi/180)\times N)$
@@ 2 Mapping Networks
Each CNN layer is followed by a batch normalization layer and ReLU activation both in multi-branch and central branch modules. The last layer of central branch is activated using $tanh$ to ensure the Super UVGS doesn’t take any ambiguous value resulting in gradient explosion or undesired artifacts. The obtained Super UVGS $S$ representation squeezes all the 3DGS attributes to a 3 dimensional image while also maintaining local and global structural correspondence among them.
@@
where $\phi_{l}(x)$ and $\phi_{l}(y)$ are feature maps extracted from pretrained layers of AlexNet[62].
Mapping Training Details: Before training the models, we normalized the different attributes in UVGS to [$-1,1$] using the same normalization functions as used in 3DGS paper[18]. The normalized UVGS maps are used to train the multi-branch forward and reverse mapping networks using MSE and LPIPS loss. We trained the mapping networks on $8\times A100~{}(80GB)$ GPUs with a Batch Size of 96 for 120 hours using Adam optimizer with a learning rate of $56e-$ 5 and set $\beta_{1}=0.5$ and $\beta_{2}=0.9$ with weight decay of $0.01$. We set the $\lambda$ for LPIPS loss to be 00 for the first 24 hours of training and gradually increased it from 1 to 10 for the remaining training in a step of 1.
@@ 2.1 Interpolation with UVGS
Figure 4: Here we show more comparison of unconditional 3D asset generation on the cars category with SOTA methods. Figure shows that DiffTF [3] produces low-quality, low-resolution cars lacking detail. While Get3D [13] and GaussianCube [59] achieve higher resolution, they suffer from 3D inconsistency, numerous artifacts, and lack richness in 3D detail. In contrast, our method generates high-quality, high-resolution objects that are 3D consistent with sharp, well-defined edges. The top three rows show the unconditional generation results of our method using ShapeNet dataset, while the bottom 3 show from Objaverse dataset.
@@ 3 LDM - Unconditional and Conditional Generation
Caption Generation To generate the relevant text captions for the objects in our dataset for conditional generation, we leverage CLIP [36], BLIP2 [20], and GPT4 [1] very similar to [24]. Specifically, we use BLIP2 to generate $N$ different captions for randomly selected 20 views from the 88 rendered views for each object in the dataset. CLIP encoders are used to encode and calculate the cosine similarity between the $N$ generated caption per view and the corresponding 20 views. The caption with max similarity is assigned to that particular view, resulting in 20 different captions for the same object. We now use GPT4 to extract a single caption distilling all the given 20 descriptions. We found that the resulting captions were very appropriate to the input objects, and thus we directly used them for conditional generation.
LDMs [40, 16] use pretrained VAEs [11] to convert the original image $x\in R^{H\times W\times 3}$ into a compact latent representation $z\in R^{h\times w\times c}$, where the forward and reverse diffusion processes are applied [40]. The VAE decoder then converts the compact latent representation back to pixels. The objective function in latent diffusion model can be written as:
$\mathbb{L}_{LDM}:=\mathbb{E}_{\epsilon(x),\epsilon\sim\mathbb{N}(0,1),t}[||\epsilon-\epsilon_{\theta}(z_{t},t)||_{2}^{2}]$ (7)
@@
Training was done using AdamW optimizer with a learning rate of $41e-$ 4 for 75 epochs on $8\times A100~{}(80GB)$ GPUs.
@@
To allow generation of objects from text, we also trained a conditional LDM, where we use Stable Diffusion (SD) [40] pipeline as it can use text prompt conditioning to guide the image generation through cross-attention. Similar to unconditional LDM, we use pretrained SD’s VAE for mapping the Super UVGS image to a latent space and back to the reconstructed Super UVGS. The text prompts are given to a pretrained CLIP [36] text encoder to generate a text embedding $c_{t}\in\mathbb{R}^{77\times 768}$, which is then passed to the UNet encoder of SD for cross-attention. We used a set of CLIP encoder and BLIP2 [20], and GPT4 [1] to generate captions for our dataset. The overall objective function for conditional LDM now becomes:
$\mathbb{L}^{C}_{LDM}:=\mathbb{E}_{\epsilon(x),\epsilon\sim\mathbb{N}(0,1),t,c_{t}}[||\epsilon-\epsilon_{\theta}(z_{t},t,c_{t})||_{2}^{2}]$ (8)
where, $\epsilon_{\theta}(\cdot,t)$ is a time-conditional U-Net [41] model, $\mathbb{N}(0,1)$ is the Normal distribution, $z_{t}$ is the latent code, and $c_{t}$ is the text embedding. Training was done using AdamW optimizer with a learning rate of $41e-$ 4 for 50 epochs on $8\times A100~{}(80GB)$ GPUs. Once trained, this conditional LDM can use used to generate text-conditioned Super UVGS images, which can later be mapped to high-quality 3DGS objects.
Figure 5: 3DGS Interpolation.
@@
Get3D [13] $68.95$ $4.38$ DreamGaussian [47] $28.51$
DiffTF [32] $89.14$ $8.73$ Shap. E [7] $30.76$
EG3D [4] $74.51$ $6.62$ LGM [49] $31.24$
GaussianCube [59] $48.44$ $4.89$ GaussianCube [59] $29.83$
@@
UVGS-S (Ours) 32.87 3.73 - -
Figure 6: Conditional Generation (more results are presented in the accompanying video).
@@ 4 Comparison with Baselines
We compare the generational capabilities of our method against various conditional and unconditional SOTA 3D object generation method. Specifically, we used the methods using multiview rendering for optimization, like DiffTF [3] and Get3D [13]. We also compared our approach again the current SOTA methods trying to give structural representation to Gaussians, including GaussianCube [59] and TriplaneGaussian [67]. We also compared against general purpose SOTA large 3D content generation models like DreamGaussian [47], LGM [49], and EG3D [4].
To compare the quality of our generation results, as a standard practice, we use FID and KID for unconditional generation, and Clip Score for text-conditioned generation. Table 2 quantitatively compares the unconditional and conditional generation results of our method again various SOTA methods. We also compare the results of our fine-tuned ShapeNet cars generation framework and the results are presented in Table 4 as UVGS-S. From these tables, it can be seen that our method performs a good job in unconditional generation of good quality 3D assets. The main reason behind this is the learned Super UVGS representation which not only maintains the appearance of the 3DGS object, but also serves as a proxy for geometrical shape by encoding all the 3DGS attributes into the same coherent feature space. Table 4 compares the CLIP Score of our text-conditioned generation results and the current SOTA methods. The unconditional and conditional qualitative comparison results are presented in Fig. 6 and Fig. 4, respectively.

### v4

@@ 5 Conclusion
ACKNOWLEDGEMENTS
A part of this work was supported by NSF CAREER grant 2143576 and ONR DURIP grant N00014-23-1-2804.
[…]
@@ 1 Spherical Mapping
Spherical Mapping: Spherical mapping [sphericalmapping2006] is a fundamental technique in computer graphics that is used to project 3D meshes onto a 2D map generally for texture mapping, where a 2D image is wrapped around a 3D object, such as a cylinder or a sphere. However, cylindrical mapping fails to capture the top and bottom parts of the object in the same UV map, and can introduce distortions for objects that extend far in the Z-direction. Hence we opted for spherical mapping the process of which involves converting 3D Cartesian coordinates ($x,y,z$) into spherical coordinates ($\rho,\theta,\phi$) and then mapping these onto a 2D plane. Algorithm [1] explains spherical unwrapping in detail for a single layer(K=1). The same process can be repeated for multiple layers, by keeping a track of opacity values.
@@
[…]
@@
[…]
@@ 2 Mapping Networks
Each CNN layer is followed by a batch normalization layer and ReLU activation both in multi-branch and central branch modules. The last layer of central branch is activated using tanhtanh to ensure the Super UVGS doesn’t take any ambiguous value resulting in gradient explosion or undesired artifacts. The obtained Super UVGS $S$ representation squeezes all the 3DGS attributes to a 3 dimensional image while also maintaining local and global structural correspondence among them.
@@
where $\phi_{l}(x)$ and $\phi_{l}(y)$ are feature maps extracted from pretrained layers of AlexNet[lpips2018].
Mapping Training Details: Before training the models, we normalized the different attributes in UVGS to [1,1-1,1] using the same normalization functions as used in 3DGS paper[3dgs2023]. The normalized UVGS maps are used to train the multi-branch forward and reverse mapping networks using MSE and LPIPS loss. We trained the mapping networks on $8\times A100~(80GB)$ GPUs with a Batch Size of 96 for 120 hours using Adam optimizer with a learning rate of 6e56e-5 and set $\beta_{1}=0.5$ and $\beta_{2}=0.9$ with weight decay of $0.01$. We set the $\lambda$ for LPIPS loss to be 0 for the first 24 hours of training and gradually increased it from 1 to 10 for the remaining training in a step of 1.
@@ 3 LDM - Unconditional and Conditional Generation
Caption Generation To generate the relevant text captions for the objects in our dataset for conditional generation, we leverage CLIP [clip], BLIP2 [blip22023], and GPT4 [gpt42023] very similar to [cap3d2024]. Specifically, we use BLIP2 to generate $N$ different captions for randomly selected 20 views from the 88 rendered views for each object in the dataset. CLIP encoders are used to encode and calculate the cosine similarity between the $N$ generated caption per view and the corresponding 20 views. The caption with max similarity is assigned to that particular view, resulting in 20 different captions for the same object. We now use GPT4 to extract a single caption distilling all the given 20 descriptions. We found that the resulting captions were very appropriate to the input objects, and thus we directly used them for conditional generation.
Figure 5: Linear interpolation between two 3DGS objects using SuperUVGS representation.
LDMs [stable_diffusion, ddim] use pretrained VAEs [transformer_vae] to convert the original image $x\in R^{H\times W\times 3}$ into a compact latent representation $z\in R^{h\times w\times c}$, where the forward and reverse diffusion processes are applied [stable_diffusion]. The VAE decoder then converts the compact latent representation back to pixels. The objective function in latent diffusion model can be written as:
[…]
@@
Training was done using AdamW optimizer with a learning rate of 1e41e-4 for 75 epochs on $8\times A100~(80GB)$ GPUs.
@@
To allow generation of objects from text, we also trained a conditional LDM, where we use Stable Diffusion (SD) [stable_diffusion] pipeline as it can use text prompt conditioning to guide the image generation through cross-attention. Similar to unconditional LDM, we use pretrained SD’s VAE for mapping the Super UVGS image to a latent space and back to the reconstructed Super UVGS. The text prompts are given to a pretrained CLIP [clip] text encoder to generate a text embedding $c_{t}\in\mathbb{R}^{77\times 768}$, which is then passed to the UNet encoder of SD for cross-attention. We used a set of CLIP encoder and BLIP2 [blip22023], and GPT4 [gpt42023] to generate captions for our dataset. The overall objective function for conditional LDM now becomes:
[…]
where, $\epsilon_{\theta}(\cdot,t)$ is a time-conditional U-Net [unet] model, $\mathbb{N}(0,1)$ is the Normal distribution, $z_{t}$ is the latent code, and $c_{t}$ is the text embedding. Training was done using AdamW optimizer with a learning rate of 1e41e-4 for 50 epochs on $8\times A100~(80GB)$ GPUs. Once trained, this conditional LDM can use used to generate text-conditioned Super UVGS images, which can later be mapped to high-quality 3DGS objects.
@@
Get3D [get3d2022] $53.17$ $4.19$ DreamGaussian [dreamgaussian2023] $28.51$
DiffTF [difftf2023] $84.57$ $8.73$ Shap. E [shapeditor2024] $30.53$
EG3D [eg3d2022] $74.51$ $6.62$ LGM [lgm2025] $30.74$
GaussianCube $34.67$ $3.72$ GaussianCube [gaussiancube2024] $30.34$
@@ 4 Comparison with Baselines
We compare the generational capabilities of our method against various conditional and unconditional SOTA 3D object generation method on ShapeNet-cars dataset. Specifically, we used the methods using multiview rendering for optimization, like DiffTF [difftf2023] and Get3D [get3d2022]. We also compared our approach again the current SOTA methods trying to give structural representation to Gaussians, including GaussianCube [gaussiancube2024] and TriplaneGaussian [triplanemeetsgs2024]. We also compared against general purpose SOTA large 3D content generation models like DreamGaussian [dreamgaussian2023], LGM [lgm2025], and EG3D [eg3d2022].
To compare the quality of our generation results, as a standard practice, we use FID and KID for unconditional generation, and Clip Score for text-conditioned generation. Table 2 quantitatively compares the unconditional and conditional generation results of our method again various SOTA methods. From this table, it can be seen that our method performs a good job in unconditional generation of good quality 3D assets. The main reason behind this is the learned Super UVGS representation which not only maintains the appearance of the 3DGS object, but also serves as a proxy for geometrical shape by encoding all the 3DGS attributes into the same coherent feature space. Table 4 compares the CLIP Score of our text-conditioned generation results and the current SOTA methods. The unconditional and conditional qualitative comparison results are presented in Fig. 7 and Fig. 6, respectively.
Figure 6: Here we show more comparison of unconditional 3D asset generation on the cars category with SOTA methods. Figure shows that DiffTF [difftf2023] produces low-quality, low-resolution cars lacking detail. While Get3D [get3d2022] achieve higher resolution, it suffers from 3D inconsistency, numerous artifacts, and lack richness in 3D detail. Similar issues are found in GaussianCube [gaussiancube2024] along with symmetric inconsistency in the results. In contrast, our method generates high-quality, high-resolution objects that are 3D consistent with sharp, well-defined edges. The top three rows show the unconditional generation results of our method using ShapeNet dataset, while the bottom 3 show from Objaverse dataset.
Figure 7: Text-conditioned generation results on various baselines and the proposed method. Our method not only generates high-quality assets for simpler objects, but also for complicated objects with intricate geometries like the wheel or the airplane.